What would be the result if you combine 33 grams of methane with 289 grams of oxygen


1
Expert's answer
2021-02-16T07:47:00-0500

CH4+2O2=CO2+2H2O


n(CH4)=m/M=33/16=2.062mol

n(O2)=m/M=289/32=9.03mol


n(CO2)=2.062mol

m(CO2)=n*M=2.062*44=90g


n(H2O)=2*2.062=4.124mol

m(H2O)=n*M=4.124*18=74g